Can you help me understand how to solve this word problem..."the ones digit of a two digit number is 5 less than the tens digit, and the number is 55 more than three times thesum of the digits. What is the number?". I don't want the answer, but wantto learn how to solve it thanks
Answer by ankor@dixie-net.com(22740)

Let x = the 10's digit
Let y = the units (ones)
then
10x+y = "the number"
:
Write an equation for each statement:
:
"the ones digit of a two digit number is 5 less than the tens digit,"
y = x - 5
:
"the number is 55 more than three times the sum of the digits."
10x + y = 55 + 3(x+y)
Remove the brackets
10x + y = 55 + 3x + 3y
Combine like terms, x on the left, y on the right
10x - 3x = 3y - y + 55
7x = 2y + 55
:
from the 1st equation we know y = (x-5), substitute for y in the above equation
find x, then find y
7x = 2(x-5) + 55
:
You should be able so solve it now, did this make sense to you?




What is the number?"